please find below a short summary of the activities related to the AO system.
People at LBTO:
The analysis of log files revealed a bug in the high order modes offload; the automatic correction has therefore been temporarily by-passed. Astigmatism correction, on the opposite, is still implemented. More in details, the HO modes are corrected by mean of a modes matrix that is averaged on a 20s sampling time: we found that averaging starts automatically before the shell has reached its final position during set-up. Hence the averaging mixes data that are sistematically inhomogeneous, giving rise to a wrog correction command.
The dome was opened at 2:00UT and the shell was set. The new astigmatism lookup table has been implemented and verified. In particular, the entity of the correction has found be more effectively than the previously calibrated LUT. Tests are needed (according to schedule) to verify its reproducibility on a inter-night time scale, taking into account that a point-to-point spread of approx 1umRMS was found on tens of minutes time scale.
The system behaved properly with the exception of two stops: from 4:25UT to 5:16UT and from 10:08UT to 10:18UT (the dome was closed at approx 11:30UT). Acts #129 and 159 were excluded from the currents control to allow operations to proceed. Showing the same behavior recorded on march 30th, faults are "compressed" into well localized series of errors: outside these stops, the mirrors keeps its shape correctly for a long time.
The gap inspection (during the rip-time) revealed that the "ear" on the shell edge appeared (gap ranging from 24 to 60um) and then came back.
The wind speed was relatively high (17m/s measured from the balcony and gusts up to 8m/s in the dome). TSS was automatically enabled when the wind speed was faster than 8m/s
A more detailed log may be found here: TelLogLBT672a20100404
- 06 Apr 2010